OpenStudy (anonymous):

Andre is making mini soft pretzels for his study group, and while he is baking, he likes to eat some that are fresh from the oven. The table shows the number of soft pretzels y he must make based on the number of students x that will attend. Use the function rule to complete the table.

OpenStudy (anonymous):

i dont understand function

OpenStudy (anonymous):

30,38,46

OpenStudy (anonymous):

The Answer is D

OpenStudy (anonymous):

30,38,46

OpenStudy (anonymous):

Which is D

OpenStudy (anonymous):

whats the rule

OpenStudy (anonymous):

add 8

OpenStudy (anonymous):

could have sworn i just answered this rule is \[f(x)=4x+6\]

OpenStudy (anonymous):

You have to add 8

OpenStudy (anonymous):

Set it up like a line, with y=mx+b. Find slope first- change in y divided by change in x 14-6 8 ---- = -- = 4 2-0 2 y=4x+b

OpenStudy (anonymous):

plug in x and y values from a point and solve for b 6=4(0)+b 6=b ur equation is y=4x+6

OpenStudy (anonymous):

So for the question mark problems, put the x in your equation and solve for y.